:1-40 Tax Rates. Based on the amounts of taxable income below, compute the federal income tax payable in 2018 on each amount assuming the taxpayers are married filing a joint return. Also, for each amount of taxable income, compute the average tax rate and the marginal tax rate a. Taxable income of $30,000. b. Taxable income of $100,000 c. Taxable income of $375,000 d. Taxable income of $700,000.
Answer #1

Solution:-

a. Taxable income of $30,000:-

Federal income tax payable in 2018 -

3,218.88

Explanation:-

$1,905 + (30,000 - 19,051) * 12%

= 3,218.88

Average tax rate:-

3,218.88 / 30,000 * 100 = 10.73%

Marginal tax rate:-

12%

b. Taxable income of $100,000:-

Federal income tax payable in 2018 -

13,876

Explanation:-

$8,904 + (100,000 - 77,400) * 22%

= 13,876

Average tax rate:-

13,876 / 100,000 * 100 = 13.88%

Marginal tax rate:-

22%

c. Taxable income of $375,000:-

Federal income tax payable in 2018 -

83,379

Explanation:-

$64,179 + (375,000 - 315,000) * 32%

= 83,379

Average tax rate:-

83,379 / 375,000 * 100 = 22.23%

Marginal tax rate:-

32%

d. Taxable income of $700,000:-

Federal income tax payable in 2018 -

198,379

Explanation:-

$161,379 + (700,000 - 600,000) * 37%

= 198,379

Average tax rate:-

198,379 / 700,000 * 100 = 28.34%

Marginal tax rate:-

37%

Please Rate or comment if you have any doubt regarding this solution